radomes.openskiesproject.org
lm7ch3nyforcuten
lm7ch3nyforcuten lm7ch3nyforcuten
  • 08-06-2016
  • Mathematics
Answered

You mix the letters M, A, T, H, E, M, A, T, I, C, A, and L thoroughly. Without looking, you draw one letter. Find the probability P(A). Write the probability as:

Aa fraction in simplest form
Ba decimal
Ca percent

Verified Answer :

Hagrid
Hagrid Hagrid
  • 17-06-2016

There are 12 letters in the mix and there are 3 letter A's. Therefore, the probability of drawing A of P(A) is:

a. a fraction in simplest form = 3/12 = 1/4
b. a decimal = 1/4 = 0.25
c. a percent = 0.25 *100% = 25%

Answer Link

Related Questions

Simplify. Express with positive exponents only: $\left(3 s^7 t^{-5}\right)^2\left(4 s^0 t\right)$
What is the definition of indemnity
What is sustainability
654 divided by some number is 218
MrJose divided his money in the ratio of 52 between Brian and CharlieBrain got the smaller amount of 156How much did Charlie receive PLEASE ANSWER FAST thank yo
How does federalism extend opportunities for citizen participation
Which lines in the raven show that the speaker has lost hope of ever being able to moce on and recover from the pain of losing lenore
What is 57 as a ratio
whats 274 in simplest form
the cell cycle is a repeating sequence of growth and during the life of a cell